solve by substitution and show work
10x + 11y = -26.5,
20x + 5y = -27.5

Answer by mananth(16949)

10x+11y =-26.5 .............1
5y=-27.5-20x ---------2
/ 5 =
y=-5.5-4x
Substitute y in (1)
10x+11( -5.5-4x)=-26.5
10 x -60.5 + -44 x = -26.5
10 x -44 x = -26.5 60.5
-34 x = 34
/ -34
x= -1
Plug the value of x in (1)
10 x + 11 y = -26.5
10 * -1 + 11 y = -26.5
-10 + 11 y = -26.5
11 y = -16.5
/ 11
y= -1.5
m.ananth@hotmail.ca